OpenStudy (i_need_help_please):

Will Fan and Medal AC=BC, AB=6 How long are each of the Legs?

OpenStudy (i_need_help_please):



hartnn (hartnn):

is C a right angle? if yes, you can use pythagoras theorem :)

OpenStudy (i_need_help_please):

yes it is.

OpenStudy (i_need_help_please):

i was trying to use the Pythagorean theorem but it wasn't working out for me. I couldn't figure out how to do it correctly with this problem.

hartnn (hartnn):

did you get something like \(x^2 +x^2 = 6^2\) ?

OpenStudy (i_need_help_please):

yes

hartnn (hartnn):

\(2x^2 = 36 \\x^2 = 18\) just take square root on both sides

OpenStudy (i_need_help_please):

sqrt18= 3sqrt 2 sqrt 36=6

hartnn (hartnn):

\(x^2 = 18 \\ x = \pm 3 \sqrt 2\) thats it!

OpenStudy (i_need_help_please):

thats what i was doing but then is 3 sqrt. 2 the lengths of the two sides because if so they dont add up to 6.

hartnn (hartnn):

why would the lengths of 2 sides add up to 6 ???

OpenStudy (i_need_help_please):

the length squared* dont add up to 6^2

OpenStudy (i_need_help_please):

to do the Pythagorean theorym

hartnn (hartnn):

is it? \((3\sqrt2)^2 + (3\sqrt 2 )^2 \\ = 18 +18 = 36 = 6^2\)

OpenStudy (i_need_help_please):

okay. so both of the lengths are definitely 3sqrt. 2?

hartnn (hartnn):

yup :)

OpenStudy (i_need_help_please):

Okay, Thanks so much!

hartnn (hartnn):

welcome ^_^